# MPLS Network Actions: Technological Overview and P4-Based Implementation on a High-Speed Switching ASIC

### v2

@@ Abstract
Index Terms:
Alternate-Marking Method, Data Plane Programming, IETF, MPLS Network Actions, Multiprotocol Label Switching, Network Slicing, P4, Wired CommunicationsI Introduction
The Multiprotocol Label Switching (MPLS) protocol defined in RFC 3031[1] has become a predominant wide-area networking technology. In MPLS, packets are equipped with labels adding domain-specific forwarding information. Packets are switched to intermediate nodes based on the assigned label. Since the standardization in 2001, the protocol has continuously evolved. In the last years, new applications have emerged that require labels to not only contain forwarding information but also information on how to process a packet. Therefore, special purpose labels (SPLs) have been defined. Examples of applications using SPLs include entropy labels which support equal-cost multi-path (ECMP) [2], and labels for No Further Fast Reroute (NFFRR) [3] which avoid the looping of packets during network failures. Further, labels can embed service function chaining (SFC) information[4, 5], and data related to operations, administration and maintenance (OAM) [6, 7].
@@
The contribution of this paper is manifold. We provide an introduction to the current state of the MNA framework and the first use cases identified by the working group. We compare the concept of the MNA framework to IPv6 EH and explore the feasibility for hardware implementations. Then, we implement the MNA framework based on the IETF WG proposal [14] in P4 on the Intel Tofino™ 2 switching ASIC using ISD. Our implementation is the first and to date the only implementation of the MNA framework. We evaluate the scalability and performance of the P4-MNA implementation. Further, we implement and evaluate two example network actions, namely the alternate-marking method (AMM) for performance measurement, and bandwidth reservation using network slicing. Based on our implementation experience, we propose to extend the signaling of hardware capabilities within an MNA-capable MPLS domain so that MPLS nodes with less capabilities can be integrated. Finally, we identify challenges arising from using ISD and propose solutions to address them in future work.
The rest of the paper is structured as follows. In Section II, we provide a primer on the MNA framework, including information on MPLS, SR-MPLS, SPLs, and MNA. An overview of identified use cases by the MPLS WG and their mechanisms is given in Section III and in Section IV, we compare IPv6 EH with MNA and review related work. In Section V, we explain the concept of the programming language P4. The implementation of the MNA framework on the Intel Tofino™ 2 is described in Section VI. In Section VII we evaluate the P4-MNA implementation and the exemplary network actions. In Section VIII, we describe constraints for header stacking and propose a signaling extension, and in Section IX, we identify challenges with mutable data. Finally, we conclude the paper in Section X.
@@ II A Primer on MPLS Network Actions (MNA)
In this section, we give a brief overview of traditional MPLS networks including forwarding in MPLS and special-purpose labels. Then, we explain the concept of the MPLS Network Actions (MNA) framework as being standardized by the IETF MPLS working group[15]. This includes the proposed MNA header encoding, and the placement of network actions in the MPLS label stack.
@@ II-A1 MPLS Forwarding
About 10 years ago, segment routing (SR) has been introduced which turned MPLS into a connection-less and source-routed packet switching technology. Labels are interpreted as segment identifiers (SIDs) which denote next intermediate nodes to where the packet is forwarded on the way to the egress LER. The label used for forwarding at an LSR is called active segment [16]. When the packet has reached that intermediate node, the corresponding label is popped. By stacking multiple labels, a source route is defined. Thus, with SR, the ingress LER pushes a label stack and intermediate hops pop these labels. There is no connection concept with SR-MPLS and therefore no signaling protocol is required to set up connections. However, SIDs need to be signaled and label stacks need to be computed. An exemplary network using SR-MPLS is shown in Figure 1(b).
@@ II-B2 The MNA Header Encoding
For network actions and AD LSEs, the encoding of an MPLS LSE is repurposed as shown in Figure 5.
@@
Optionally, subsequent opcode LSEs (Format C) or AD LSEs (Format D) follow. The Format C LSE is a simplified version of the Format B encoding. The Format D LSE contains AD which relates to the preceding Format B, or Format C LSE. AD can be carried either as part of a Format B or Format C LSE, or as a separate AD LSE in Format D. In the Format C LSE, 20 bits are available for opcode-specific AD. Finally, the Format D LSE contains bits for opcode-specific AD. The semantics of the AD are left to the predefined network action. A NAS can contain up to 16 network actions and AD LSEs. Together with the NAS indicator, a NAS therefore contains up to 17 LSEs.
@@ II-B3 MNA Scopes and MNA with SR-MPLS
The MPLS stack in Figure 6 encodes MPLS forwarding labels for three segments. In this example, LSRs parse the entire MPLS stack. Therefore, in Figure 6, executes network actions in the select-scoped NAS located below the top-of-stack forwarding label, and the HBH-scoped NAS located below the forwarding label . pops the select-scoped NAS after popping . executes network actions in the HBH-scoped NAS. executes network actions in the HBH-scoped NAS, the select-scoped NAS, and the I2E-scoped NAS located below its forwarding label.
@@ III Identified Use Cases For the MNA Framework
In this section, we describe five use cases identified by the IETF working group[19], namely no further fast reroute (NFFRR), in-situ OAM, SFC, AMM, and network slicing. Many of the identified use cases for MNA exist as a technology-agnostic mechanism and are implemented and adapted for the MPLS protocol. Table I shows a summary of the identified use cases for the MNA framework. For each use case, the technology-agnostic mechanism, the adaption to MPLS, and a proposal for the MNA framework are shown. In the following sections, the use cases are described in more detail.
@@ IV-A Comparison of IPv6 Extension Headers with MNA
RFC 2460[37] defines IPv6 EHs to expand the functionality of IPv6. IPv6 EHs contain options that indicate special operations such as fragmentation or allow for segment routing (SR) over IPv6 (SRv6). IPv6 allows for global, inter-domain communication between devices. This makes extensions, e.g., new options in an EH, difficult to deploy because devices worldwide must implement new extensions to take advantage of them. In contrast, MPLS is typically deployed by providers as a Layer 2 and Layer 3 technique within their own domain. This facilitates the deployment of extensions, such as the MNA framework or new network actions in the MNA framework. The usefulness of MNA in a providers domain does not depend on MNA deployment in other domains.
@@ VII-B1 Testbed and Methodology
In Figure 20, packets are sent to an ingress port of the Intel Tofino™ 2 switch. Each ingress port of the P4-MNA switch represents an LSR processing network actions. After processing the network actions in the stack, LSRs pop the top-of-stack label and forward packets according to the active segment. For label , , and packets are sent to a different ingress port on the same Intel Tofino™ 2 switch. LSR forwards packets back to the traffic generator P4TG for traffic analysis and measurement. In the testbed setup, all LSRs operate on the same physical device and share a common time reference.
@@ VII-B2 Performance Results
This evaluation has two goals. First, we want to verify if P4-MNA is capable of achieving line rate processing of network actions. Second, we analyze the impact on the processing delay resulting from network actions in the MPLS stack. We measure the packet loss PL and the RTT of the returned traffic using the measurement capabilities of the traffic generator P4TG. The mean results of the RTT and packet loss measurement are shown in Table III. For clarity, we omit confidence intervals as their width was less than of the measured average for a confidence level of .
@@
Table III shows that no packet loss was recorded, i.e., P4-MNA achieves line rate forwarding of with MPLS traffic in the experiments E1 – E4. Adding network actions increases the RTT marginally by from in E1 to in E2 – E4. A packet is processed once by each LSR, i.e., three times in total222The per-packet processing delay is constant on the Intel Tofino™.. The latency added by one LSR in E2 – E4 is therefore .
@@ VII-C2 Validation of Link-Specific Packet Loss
The mean results from experiment E6 are shown in Table IV. The width of the confidence intervals was less than of the measured average and therefore, they are not shown.
@@
Table IV shows that the calculated link-specific packet loss rate matches the configured drop probabilities for each LSR. For LSR , , for LSR , and for LSR packet loss was measured. This matches the configured drop probability at each LSR exactly. Thus, the implemented AMM network action is capable of precisely measuring link-specific packet loss.
@@ VII-D2 Validation of Bandwidth Reservations
First, we measure the packet loss per stream and the RTT without generating interference traffic, i.e., no congestion is caused on the link. Next, we add interference traffic but do not apply the network slicing action, leading to congestion. Finally, we apply the network slicing action. The mean measured packet loss per stream for is shown in Figure 23. The width of the confidence intervals was less than of the measured average and therefore are invisible.
@@ VIII-A Constraints for Header Stacking due to Readable Label Depth (RLD)
MPLS nodes have a limited depth of LSEs they can parse without performance impact resulting from their available resources. This value is called readable label depth (RLD)[8]. The RLD is a critical parameter of LSRs when deploying MNA in MPLS networks. In the current proposal, LSRs signal their RLD to the ingress LERs using a routing protocol such as IS-IS [54] or OSPF [55]. The ingress LER must ensure that a NAS determined for a node is within the RLD when the packets reach that node. This is done based on the signaled RLD parameter. An HBH-scoped NAS can be located at the bottom of stack. However, if this HBH-scoped NAS is not within the RLD of an intermediate node, the ingress LER places copies of that NAS within the RLD to ensure that the NAS can be found by that node. Only the topmost HBH NAS copy is processed by a node. An example is given in Figure 24 and further described below.
@@ VIII-B Constraints for Header Stacking due to Maximum NAS Sizes
This holds under the assumption that the hardware implementation cannot share the memory for parsing the select-scoped NAS, the HBH-scoped NAS, and the in-between stack, which is the case for P4-MNA. The relation between the three arrays and the RLD is depicted on the left side in Figure 25. In case of P4-MNA, , therefore, the in-between stack can contain up to 16 LSEs. They contain any LSE between the two NAS, i.e., forwarding labels, special purpose labels, and network actions. This seems a low value compared to a maximum NAS size. Moreover, we conclude from Equation 6 that MNA-capable hardware must have an RLD of at least 34 LSEs, which may be problematic for some hardware platforms.
@@ IX Challenges with Mutable Data in an ISD Implementation
X Conclusion
@@
- SR
- segment routing
- SID
- segment identifier
@@
- OAM
- operations, administration and maintenance
- SFC
- service function chaining
- ISD
- in-stack data
- PSD
- post-stack data
- MPLS
- Multiprotocol Label Switching
- MNA
- MPLS Network Actions
- AD
- ancillary data
- NAS
- Network Action Sub-stack
- LSP
- Label Switched Path
- LSE
- Label Stack Entry
- LSR
- Label Switching Router
- LER
- Label Edge Router
@@
- NASL
- Network Action Sub-stack Length
- NAL
- Network Action Length
- FRR
- Fast Reroute
- NSH
- network service header
- AMM
- alternate-marking method
- NRP
- network resource partition
- P4
- Programming Protocol-independent Packet Processors
- MAT
- match+action table
- CBR
- constant bit rate

### v3

@@ Abstract
Alternate-Marking Method, Data Plane Programming, IETF, MPLS Network Actions, Multiprotocol Label Switching, Network Slicing, P4, Wired Communications
I Introduction
[lines=2]T he Multiprotocol Label Switching (MPLS) protocol defined in RFC 3031 [3] has become a predominant wide-area networking technology. In MPLS, packets are equipped with labels adding domain-specific forwarding information. Packets are switched to intermediate nodes based on the assigned label. Since the standardization in 2001, the protocol has continuously evolved. In the last years, new applications have emerged that require labels to not only contain forwarding information but also information on how to process a packet. Therefore, special purpose labels (SPLs) have been defined. Examples of applications using SPLs include entropy labels which support equal-cost multi-path (ECMP) [4], and labels for No Further Fast Reroute (NFFRR) [5] which avoid the looping of packets during network failures. Further, labels can embed service function chaining (SFC) information[6, 7], and data related to operations, administration and maintenance (OAM) [8, 9].
@@ I Introduction
The contribution of this paper is manifold. We provide an introduction to the current state of the MNA framework and the first use cases identified by the working group. We compare the concept of the MNA framework to IPv6 EH and explore the feasibility for hardware implementations. Then, we implement the MNA framework based on the IETF MPLS working group (WG) proposal [16] in P4 on the Intel Tofino™ 2 switching ASIC using ISD. Our implementation is the first and to date the only implementation of the MNA framework. We evaluate the scalability and performance of the P4-MNA implementation. Further, we implement and evaluate two example network actions, namely the alternate-marking method (AMM) for performance measurement, and bandwidth reservation using network slicing. Based on our implementation experience, we propose to extend the signaling of hardware capabilities within an MNA-capable MPLS domain so that MPLS nodes with less capabilities can be integrated. Finally, we identify challenges arising from using ISD, propose solutions to address them in future work, and summarize security considerations of the MNA framework.
The rest of the paper is structured as follows. In Section II, we provide a primer on the MNA framework, including information on MPLS, SR-MPLS, SPLs, and MNA. An overview of identified use cases by the IETF MPLS WG and their mechanisms is given in Section III and in Section IV, we compare IPv6 EH with MNA and review related work. In Section V, we explain the concept of the programming language P4. The implementation of the MNA framework on the Intel Tofino™ 2 is described in Section VI. In Section VII we evaluate the P4-MNA implementation and the exemplary network actions. In Section VIII, we describe constraints for header stacking and propose a signaling extension, and in Section IX, we identify challenges with mutable data. In Section X, we describe security risks currently considered in the MNA framework. Finally, we conclude the paper in Section XI.
@@ II A Primer on MPLS Network Actions (MNA)
In this section, we give a brief overview of traditional MPLS networks including forwarding in MPLS and special-purpose labels. Then, we explain the concept of the MPLS Network Actions (MNA) framework as being standardized by the IETF MPLS WG[17]. This includes the proposed MNA header encoding, and the placement of network actions in the MPLS label stack.
@@ II-A1 MPLS Forwarding
About 10 years ago, segment routing (SR) has been introduced which turned MPLS into a connection-less and source-routed packet switching technology. Labels are interpreted as segment identifiers (SIDs) which denote next intermediate nodes to where the packet is forwarded on the way to the egress LER. In SR, there are two types of segments. An adjacency segment describes a strict forwarding instruction over a specific link between two nodes while a prefix segment describes a loose forwarding instruction to a prefix over multiple hops [18]. For an adjacency segment, a node pops the top-of-stack label and forwards the packet to the next segment [19]. For a prefix segment, the top-of-stack label is not popped until the prefix is reached. By stacking multiple labels, a source route is defined. Thus, with SR, the ingress LER pushes a label stack and intermediate hops pop these labels. There is no connection concept with SR-MPLS and therefore no signaling protocol is required to set up connections. However, SIDs need to be signaled and label stacks need to be computed. An exemplary network using SR-MPLS with adjacency segments is shown in Figure 1(b). In both examples, penultimate hop popping is not applied, i.e., the egress node receives a label, pops it, and forwards the packet based on the underlying Layer 3 information.
@@ II-B2 The MNA Header Encoding
Optionally, subsequent opcode LSEs (Format C) or AD LSEs (Format D) follow. The Format C LSE is a simplified version of the Format B encoding. The Format D LSE contains AD which relates to the preceding Format B, or Format C LSE. AD can be carried either as part of a Format B or Format C LSE, or as a separate AD LSE in Format D. In the Format C LSE, 20 bits are available for opcode-specific AD. Finally, the Format D LSE contains bits for opcode-specific AD. The semantics of the AD are left to the predefined network action. A NAS can contain up to 16 network actions and AD LSEs. Together with the NAS indicator, a NAS therefore contains up to 17 LSEs. After exposing a NAS to the top, it must be popped.
@@ II-B3 MNA Scopes and MNA with SR-MPLS
The MPLS stack in Figure 6 encodes MPLS forwarding labels for adjacency segments. In this example, LSRs parse the entire MPLS stack. Therefore, in Figure 6, executes network actions in the select-scoped NAS located below the top-of-stack forwarding label, and the HBH-scoped NAS located below the forwarding label . pops the select-scoped NAS after popping . executes network actions in the HBH-scoped NAS. executes network actions in the HBH-scoped NAS, the select-scoped NAS, and the I2E-scoped NAS located below its forwarding label. In this example, penultimate hop popping is not applied. With penultimate hop popping, removes the last forwarding label but must not remove the NAS exposed to the top. Then, may receive a NAS at the top of stack [16].
@@ III Identified Use Cases For the MNA Framework
In this section, we describe five use cases identified by the IETF working group[22], namely no further fast reroute (NFFRR), in-situ OAM, SFC, AMM, and network slicing. Many of the identified use cases for MNA exist as a technology-agnostic mechanism and are implemented and adapted for the MPLS protocol. Table 1 shows a summary of the identified use cases for the MNA framework. For each use case, the technology-agnostic mechanism, the adaption to MPLS, and a proposal for the MNA framework are shown. In the following sections, the use cases are described in more detail.
@@ IV-A Comparison of IPv6 Extension Headers with MNA
RFC 2460[40] defines IPv6 EHs to expand the functionality of IPv6. IPv6 EHs contain options that indicate special operations such as fragmentation or allow for segment routing (SR) over IPv6 (SRv6). Table 2 compares the concepts from IPv6 EH with the MNA framework.
@@
IPv6 allows for global, inter-domain communication between devices. This makes it difficult to deploy extensions, such as new options in a EH, because devices around the world must implement new extensions to take advantage of them. In contrast, MPLS is typically deployed by providers as a Layer 2 and Layer 3 technique within their own domain. This facilitates the deployment of extensions, such as the MNA framework or new network actions in the MNA framework. The usefulness of MNA in a provider’s domain does not depend on the deployment of MNA in other domains as providers can choose which network actions to support in their own domain.
@@ VII-A Scalability Analysis of P4-MNA
More complex network actions may require more instructions that may not fit in the pipeline. In this case, recirculation is required. In a recirculation, a packet is looped back to an ingress port where it is processed again, i.e., the required instructions are applied. An advantage of this is the implementation of more complex functions such as shown by [52]. However, recirculation comes at the cost of bandwidth. To avoid this, simple network actions should be preferred. The network actions currently proposed by the WG are simple enough to not require recirculation. However, a combination of many simple network actions may also exceed the packet processing resources within a single packet processing cycle and require recirculation.
@@ VII-B1 Testbed and Methodology
In Figure 20, packets are sent to an ingress port of the Intel Tofino™ 2 switch running the P4-MNA implementation. Each ingress port of the P4-MNA switch represents an LSR processing network actions. After processing the network actions in the stack, LSRs pop the top-of-stack label and forward packets accordingly. For label , , and packets are sent to a different ingress port on the same Intel Tofino™ 2 switch. LSR forwards packets back to the traffic generator P4TG for traffic analysis and measurement. In the testbed setup, all LSRs operate on the same physical device and share a common time reference.
@@ VII-B2 Performance Results
This evaluation has two goals. First, we want to verify if P4-MNA is capable of achieving line rate processing of network actions. Second, we analyze the impact on the processing delay resulting from network actions in the MPLS stack. We measure the packet loss PL and the RTT of the returned traffic using the measurement capabilities of the traffic generator P4TG. The mean results of the RTT and packet loss measurement are shown in Table 3. For clarity, we omit confidence intervals as their width was less than of the measured average for a confidence level of .
@@
Table 3 shows that no packet loss was recorded, i.e., P4-MNA achieves line rate forwarding of with MPLS traffic in the experiments E1 – E4. Adding network actions increases the RTT marginally by from in E1 to in E2 – E4. A packet is processed once by each LSR, i.e., three times in total222The per-packet processing delay is constant on the Intel Tofino™.. The latency added by one LSR in E2 – E4 is therefore .
@@ VII-C2 Validation of Link-Specific Packet Loss
The mean results from experiment E6 are shown in Table 4. The width of the confidence intervals was less than of the measured average and therefore, they are not shown.
@@
Table 4 shows that the calculated link-specific packet loss rate matches the configured drop probabilities for each LSR. For LSR , , for LSR , and for LSR packet loss was measured. This matches the configured drop probability at each LSR exactly. Thus, the implemented AMM network action is capable of precisely measuring link-specific packet loss.
@@ VII-D2 Validation of Bandwidth Reservations
First, we measure the packet loss per stream and the RTT without generating interference traffic, i.e., no congestion is caused on the link. Next, we add interference traffic but do not apply the network slicing action, leading to congestion. Finally, we apply the network slicing action. The mean measured packet loss per stream for is shown in Figure 23. The width of the confidence intervals was less than of the measured average and therefore invisible.
@@ VIII-A Constraints for Header Stacking due to Readable Label Depth (RLD)
In general, hardware can only parse limited bytes of a packet header. This limitation results from the hardware resources available on the device such as memory. Additionally, the complexity of a packet parser is limited to facilitate packet processing within a single packet cycle. Header sizes that exceed the available resources for parsing cannot be processed in a single packet cycle and, therefore, are an obstacle for packet forwarding at line rate. Thus, for practical deployment, the relevant information in the header must fit within the readable bytes. In MNA, this translates to the so-called readable label depth (RLD) [10] which indicates the number of LSEs a node can parse.
The RLD is a critical parameter of LSRs when deploying MNA in MPLS networks. In the current proposal, LSRs signal their RLD to the ingress LERs using a routing protocol such as IS-IS [58] or OSPF [59]. The ingress LER must ensure that a NAS determined for a node is within the RLD when the packets reach that node. This is done based on the signaled RLD parameter. An HBH-scoped NAS can be located at the bottom of stack. However, if this HBH-scoped NAS is not within the RLD of an intermediate node, the ingress LER places copies of that NAS within the RLD to ensure that the NAS can be found by that node. Only the topmost HBH NAS copy is processed by a node. An example is given in Figure 24 and further described below.
@@ VIII-B Constraints for Header Stacking due to Maximum NAS Sizes
This holds under the assumption that the hardware implementation cannot share the memory for parsing the select-scoped NAS, the HBH-scoped NAS, and the in-between stack, which is the case for P4-MNA. The relation between the three arrays and the RLD is depicted on the left side in Figure 25. In case of P4-MNA, , therefore, the in-between stack can contain up to 16 LSEs. They contain any LSE between the two NAS, i.e., forwarding labels, special purpose labels, and network actions. This seems a low value compared to a maximum NAS size. Moreover, we conclude from Equation 6 that MNA-capable hardware must have an RLD of at least 35 LSEs, which may be problematic for some hardware platforms.
@@ IX Challenges with Mutable Data in an ISD Implementation
X Security Considerations
In this section we describe security risks and considerations arising from the use of the MNA framework. Currently, there are three security risks that are considered in the MNA drafts [10, 16]. The first is link-level security, the second is MNA information originating from other domains, and the third is the syntax and semantics of network actions.
X-A Link-Level Security
The information contained in network actions is sensitive and may be exploited or manipulated for attacks. Thus, MNA traffic should be protected from eavesdropping and manipulation. However, the MPLS protocol does not have a built-in security mechanism. Link-level security mechanisms such as MACSec can be employed to prevent eavesdropping on traffic using confidentiality and authentication. Further, there is a draft for opportunistic encryption for hop-by-hop and end-to-end encryption in MPLS [64]. However, end-to-end encryption is not feasible with hop-by-hop MNA processing as nodes need to inspect the contents of the MPLS stack [10].
X-B MNA Information Originating from Other MPLS Domains
An MPLS network must be protected from processing MPLS labels originating outside the network, e.g., label stacks forged by an attacker [10, 16]. LSRs must process only label stacks including network actions intended for them. This is trivial within a single domain, but challenging in a multi-domain context. We consider the example in Figure 27 where traffic from customer domain is tunneled through a provider domain into customer domain . and are under the same administrative control while is not. LER pushes LSEs including network actions onto packets that are intended for processing by and but not for processing by and .
In the provider domain , the LERs and must be protected from erroneously processing the HBH-scoped NAS which is intended only for LSRs and . Several options are possible. First, if MNA is not supported in the provider domain, the network actions are not executed and can be forwarded through the provider domain. However, in this case, the provider domain cannot take advantage of MNA. Second, if the ingress LER has an RLD equal to or greater than all other LSRs in the provider domain, it can drop all incoming packets that contain network actions in reach for any LSR of . However, this does not allow MNA traffic to pass through the provider domain. In a third solution, the ingress LER pushes an entire new MPLS stack including a bottom of stack bit so that all LSRs in the provider domain stop parsing the stack beyond that bit. This also prevents them from executing the HBH-scoped network actions because it isolates the MNA information into two separate domain-specific stacks [10].
The LSRs in the customer domain must be protected from executing network actions that may be inserted by an adversarial provider. As a simple solution, the ingress LER may drop any incoming traffic containing network actions. However, then network actions cannot be used between and which are under the same administrative control. To support MNA, the ingress LER may forward external packets from a set of non-critical network actions such as NFFRR and block critical network actions such as for network slicing. As a third solution the customer trusts its service provider and accepts all MNA traffic for execution in domain .
X-C Syntax and Semantics of Network Actions
Operators must ensure that ingress LERs enforce syntactically correct NAS. Otherwise, the MPLS stack will be corrupted resulting in packet loss. In addition, network actions are semantically unbounded, i.e., they are not restricted to a fixed feature set but can describe arbitrary operations. For example, a network action could be defined that unintentionally enables manipulating a node’s memory to compromise forwarding or facilitate eavesdropping. The IETF must therefore ensure that only secure network actions are defined that do not compromise the security of the network. Some opcodes are reserved for locally defined network actions. In this case, implementers and network operators must ensure that these network actions are secure [16].
XI Conclusion
@@
Finally, we summarized the security risks currently being considered by the working group which include link-level security, MNA information originating from other MPLS domains, and the syntax and semantics of network actions.
Erratum
After publication in IEEE OJCOMS [1], it was identified that Figures 8(a), 8(b), and 27 were incorrectly rendered due to duplication and formatting issues introduced during production. An official Erratum has been published and is linked on the IEEE Xplore page [2]. The figures included in this version are correct.
@@
- AD
- ancillary data
- AMM
- alternate-marking method
- bSPL
- base Special Purpose Label
- CBR
- constant bit rate
@@
- ISD
- in-stack data
- LER
- Label Edge Router
- LSE
- Label Stack Entry
- LSP
- Label Switched Path
- LSR
- Label Switching Router
- MAT
- match+action table
- MNA
- MPLS Network Actions
- MPLS
- Multiprotocol Label Switching
- NAL
- Network Action Length
- NASL
- Network Action Sub-stack Length
- NAS
- Network Action Sub-stack
- NRP
- network resource partition
- NSH
- network service header
- OAM
- operations, administration and maintenance
- P4
- Programming Protocol-independent Packet Processors
- PSD
- post-stack data
@@
- SFC
- service function chaining
- SID
- segment identifier
- SR
- segment routing
- TC
- traffic class